# LogPrism: Unifying Structure and Variable Encoding for Effective Log Compression

### v1

@@ I Introduction
This design maximizes compression ratios by capturing deep contextual redundancies while simultaneously accelerating processing speed. By filtering out the majority of high-frequency patterns in the early stages, LogPrism drastically reduces the computational load on the final, expensive residual processing stage. We further enhance scalability through a parallel-aware architecture that supports fine-grained concurrency. We evaluate LogPrism on 16 benchmark datasets from LogHub [44], covering diverse system types and log formats. The results demonstrate that LogPrism establishes a new state-of-the-art in both effectiveness and efficiency, achieving the highest compression ratio on 13 out of 16 datasets and outperforming the leading baseline Denum by up to 25.25%, all while delivering the fastest compression speed (averaging 29.87 MB/s).
@@ II-A Parser-based Log Compression
In the realm of log compression, parser-based methods represent the predominant approach. They leverage the inherent semi-structured nature of logs, typically composed of static format strings and dynamic parameters. The process begins with a critical prerequisite step, i.e., log parsing. As illustrated in Fig. 1, parsers like Drain [14] are employed decompose raw log messages, separating structured headers (e.g., timestamps, log levels) from the free-form body, and then segment the body into an invariant event template and its corresponding variables. This structured representation enables the core compression mechanism, where repetitive template text is replaced with compact identifiers (stored only once in a dictionary) and the extracted parameters are aggregated for subsequent encoding using specialized compression techniques. The resulting output, which is a highly regular stream of template IDs and parameter arrays with significantly reduced entropy, is then processed by general-purpose algorithms like LZMA or gzip to eliminate any remaining statistical redundancy. Several representative log compressors [23, 33, 21, 38] exemplify different design philosophies within this framework.
@@ II-C A Motivating Example for Unified Redundancy Encoding
A straightforward way to implementing this idea is to build a prefix tree (i.e., a trie) over the sequences of log tokens and assign IDs to frequent paths, thereby identifying recurring log (sub)sequences. However, this naive construction presents two fundamental limitations. First, high-cardinality fields induce many branches, causing the number of nodes to grow with the product of distinct values per position. This results in prohibitive storage consumption and increased lookup latency due to structural inefficiency. Second, a trie only encodes prefixes. If an infrequent token appears early, the match stops and the rest of the log line is no longer compressible, even when the suffix is highly regular (e.g., uid=509 euid=0 after a rare rhost in L6). These constraints motivate our design of LogPrism, an effective log compressor which moves beyond pure prefix matching and supports compact, gap-tolerant aggregation of frequent token sequences.
@@ III-B Variable Subtree Encoding
The construction of the structural context tree leverages only stable tokens, resulting in a highly compact skeleton where all variable tokens are aggregated at the terminal nodes. To capture the deep correlations between the log structure and its parameters, we can conceptually extend the URT by attaching a prefix subtree to each terminal node, using the varLists of the associated log entries as input paths. By doing so, we unify the structural skeleton and variables into a single, continuous representation. However, this presents two critical challenges. First, a naive sequential insertion of variables inevitably triggers a secondary branch explosion. As illustrated in Fig. 9(a), if variables are processed in their original order (e.g., $v_{0}\to v_{1}\to v_{2}\to v_{3}$, a high-cardinality variable may appear early in the sequence (e.g., a rhost at $v_{1}$), forcing the subtree to branch at the top layers. Consequently, even if later variables form a highly frequent pattern (e.g., uid=509 $\to$ euid=0), it is fragmented into multiple separate branches, preventing efficient compression. Second, we require a robust mechanism to distinguish meaningful patterns from random noise. Since variables are high-entropy, assigning a unique ID to every combination would result in a prohibitively large dictionary and compromise the compression benefits. Therefore, we must selectively encode only frequent, co-occurring variable combinations (the “signal”) into the URT, while actively filtering out rare or random values (the “noise”) to be handled as residuals (Sec. III-C).
@@ III-B1 Variable Reordering for Efficient Subtree Construction
The optimization begins with high-frequency filtering, which prunes infrequent variables, as shown in Fig. 10(a). For each variable position within a log group, we compute the frequency of every unique value and discard any that fall below a pre-defined threshold $\alpha$ (set to 1% of the total log count in the group), e.g., the rare rhost=srv2.alfahost.nl. If all values at a given position are filtered out, the entire position is excluded from subsequent analysis. For instance, in the httpd log group (pathID=1) shown in Fig. 8, the sole variable $v_{0}$ (transaction_id) is removed entirely because both of its values (5001 and 5002) are too infrequent.
@@ III-B2 Subtree Construction and Unified Redundancy Encoding
The construction process is illustrated in Fig. 10(b). As each log’s varList is traversed according to the optimal order (i.e., $v_{3}\to v_{0}\to v_{2}\to v_{1}$), a path is extended in the subtree. Every node in this tree maintains a cnt attribute, a key metric that records the number of logs whose reordered variables share that specific prefix. For logs like L5 and L7, which are composed entirely of frequent variables, a complete path is formed, and the cnt value of every node along that path is incremented. In contrast, the traversal for log L6 terminates prematurely because its variable rhost=srv2.alfahost.nl was filtered out as a low-frequency value. Consequently, L6 only increments the cnt values for its matched prefix (euid=0 $\to$ user=test1 $\to$ uid=509). This distinction is crucial, as the cnt attribute now precisely tracks the frequency of both complete and partial patterns. Next, a pruning operation is performed, which acts as a second layer of filtering, removing paths composed of individually frequent variables but whose combination is rare. Any node whose cnt falls below a pre-defined pruning threshold $\beta$ (set to 0.5% of the total log count in the group) is removed. While no nodes are pruned in our simplified example (Fig. 10), this is crucial for eliminating noise in complex datasets.
The most critical step is the assignment of new universal pathIDs. To maximize compression efficiency, identifiers are assigned only to nodes that represent the termination of a high-frequency aggregate pattern. This ensures that every pathID corresponds to a meaningful, recurring token combination, preventing an explosion in dictionary size due to rare partial matches. We term these nodes “stable endpoints” and identify them using the cnt attribute under two conditions. First, any leaf node is inherently a stable endpoint, as it represents the explicit termination of a pattern that has already survived the high-frequency filtering process (i.e., $\beta$). For example, the rhost=pokemon1.cs.edu node in Fig. 10(b) is assigned pathID=5 because it marks the end of the pattern for logs L5 and L7. Second, a non-leaf node is designated a stable endpoint if the difference between its cnt and the sum of its children’s cnt values is not less than the threshold $\beta$. This “residual count” represents the number of log entries whose pattern matches exactly up to this node but does not continue to any of its high-frequency children. By enforcing the threshold, we ensure that we only create a new ID if a substantial number of logs terminate at this specific intermediate point. For instance, the uid=509 node has a cnt of 3, while its only child has a cnt of 2. The residual count of 1 indicates that one log (L6) terminates its match here. If this residual meets the threshold, the node is marked as a stable endpoint and assigned pathID=4, allowing LogPrism to hierarchically encode both complete patterns and frequent sub-patterns.
@@ III-C2 Just-in-Time Residual Templatization
Although this stage borrows the principle of numeric classification from Denum, our hierarchical data distillation strategy is fundamentally different. Denum performs an indiscriminate, global numeric/non-numeric split from the outset, which can prematurely destroy valuable contextual correlations. In contrast, LogPrism has already processed the vast majority of high-frequency tokens (including many numerics) via unified redundancy encoding in the first two stages. Consequently, the computationally intensive parsing logic of this stage is applied only to a minimal subset of true residual data. This targeted approach provides a dual advantage: a higher compression ratio, as it preserves the “structure + variable” correlations, and faster compression speed, by confining the expensive variable processing to a much smaller dataset.
@@ IV-B RQ1: The Compression Ratio of LogPrism
Table II presents the compression ratio comparison across all benchmarks. LogPrism achieves the highest compression ratio on 13 of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
@@
Comparison with log-specific compressors: LogPrism advances the state-of-the-art among log-specific methods. Compared to LogReducer, CR improvements reach up to 58.12%. Against LogShrink, the maximum gain is 53.31%. Notably, even compared to the current leading method Denum, LogPrism achieves substantial improvements of up to 25.25% on Linux and 16.34% on Thunderbird. The LogZip result for Thunderbird is marked as unavailable (–) because, as reported in the LogShrink paper [21], LogZip failed to complete parsing within one week.
@@
Android 7.742 18.857 12.787 19.370 25.165 20.776 21.857 32.494 32.552
Apache 21.308 25.186 29.557 31.688 30.375 43.028 55.940 58.517 64.250
BGL 12.927 17.637 15.461 18.927 32.655 38.600 42.385 41.804 46.907
Hadoop 20.485 36.095 32.598 32.110 35.008 52.830 60.091 78.546 77.915
HDFS 10.636 13.559 14.059 19.155 26.666 22.634 27.319 25.670 26.410
HealthApp 10.957 13.431 13.843 15.337 12.632 31.694 39.072 44.472 50.116
HPC 11.263 15.076 12.756 14.822 27.208 32.070 35.878 45.275 45.399
Linux 11.232 16.677 14.695 18.508 23.368 25.213 29.252 30.449 38.137
Mac 11.733 22.159 18.074 28.469 26.306 35.251 39.860 40.789 44.762
OpenSSH 16.828 18.918 22.865 31.977 42.606 86.699 103.175 101.654 112.910
OpenStack 12.158 14.437 15.231 17.429 17.258 16.701 22.157 22.238 23.267
Proxifier 15.716 18.982 23.619 25.489 21.493 25.501 27.029 27.288 29.759
Spark 17.825 19.908 26.497 30.614 20.825 59.470 59.739 59.470 52.496
Thunderbird 16.462 27.309 25.428 33.026 49.185 48.434 63.824 74.252
Windows 17.798 202.568 67.533 61.083 310.596 342.975 456.301 481.350 488.020
Zookeeper 25.979 27.667 36.156 38.931 47.373 94.562 116.981 135.251 142.313
This systematic performance advantage of LogPrism stems from a fundamental architectural shift. Traditional methods (whether the “parse-then-compress” workflows or Denum’s “numeric/non-numeric” split) perform an irreversible token categorization early in the process. This leads to the loss of contextual correlations that spans these category boundaries. In contrast, LogPrism’s unified redundancy encoding paradigm co-designs structural extraction with pattern encoding. This enables representing entire, highly correlated “structure + variable” token sequences with single path identifiers, where other methods require one template ID plus multiple independent variable IDs. Our hierarchical data distillation strategy prioritizes encoding high-value aggregate patterns while preserving full context, allowing it to exploit deep redundancies in log data.
On the three remaining datasets where LogPrism does not achieve the highest CR (Hadoop, HDFS, Spark), performance remains highly competitive. On Hadoop and HDFS, the differences from optimal are minimal (0.80% and 3.33%, respectively). The larger gap on Spark reflects the characteristics of the dataset, i.e., Spark logs are dominated by simple, repetitive templates with single continuously changing numeric values (e.g., over a million records like "Update row <*>"[38]. In such scenarios, the deep correlation mining of LogPrism’s second stage offers less advantage compared to methods specifically optimized for simple numeric streams. Nevertheless, as we demonstrate in RQ2, LogPrism achieves a better performance balance in these cases with significantly higher compression speed.
Summary for RQ1: LogPrism surpasses existing general-purpose and log-specific compressors in compression ratio on 13 out of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
@@ IV-C RQ2: The Compression Speed of LogPrism
The observed variation in LogPrism’s speed across datasets is primarily driven by dataset size, which dictates the effectiveness of chunk-level parallelization. For example, on the small Linux dataset (25,567 lines), the volume is insufficient to fill even a single 100K-line processing chunk, precluding the standard multi-threaded acceleration used by all compressors. Despite this constraint, the standard LogPrism configuration achieves the best performance (8.55 MB/s) among all baselines. Moreover, LogPrism-P provides substantial additional speedup by leveraging the co-design of hierarchical workload reduction and internal fine-grained parallelism. This ensures high efficiency even in edge cases. On the Linux dataset, LogPrism-P (10.23 MB/s) is significantly faster than Denum (4.97 MB/s) because its hierarchical strategy minimizes the computational overhead, while the internal worker threads maximize CPU utilization within the single active chunk.
@@ IV-D RQ3: Ablation Analysis of LogPrism’s Different Stages
Android 18.857 25.564 32.552
Apache 25.186 61.887 64.250
BGL 17.637 46.548 46.907
Hadoop 36.095 75.170 77.915
HDFS 13.559 24.282 26.410
HealthApp 13.431 46.932 50.116
HPC 15.076 43.880 45.399
Linux 16.677 37.356 38.137
Mac 22.159 37.373 44.762
OpenSSH 18.918 95.969 112.910
OpenStack 14.437 21.868 23.267
Proxifier 18.982 26.664 29.759
Spark 19.908 50.656 52.496
Thunderbird 27.309 59.876 74.252
Windows 202.568 490.787 488.020
Zookeeper 27.667 118.531 142.313
Compression Speed Analysis: The ablation study also reveals the critical role of Stage 2 as a performance accelerator. Despite adding another analysis step, the full LogPrism model is $25.93\%$ faster on average than the structurally simpler LogPrism (S1+S3) variant. This counter-intuitive result validates our hierarchical data distillation strategy. In the LogPrism (S1+S3) configuration, every variable identified in Stage 1 must be processed by Stage 3, which relies on computationally expensive regex templatization. In contrast, the full LogPrism model only supplies Stage 3 with a smaller set of “true long-tail residuals.” By handling the majority of high-frequency variables in the efficient Stage 2, LogPrism drastically reduces the workload of the most time-consuming part of the pipeline, thereby lowering overall computational overhead and increasing throughput.

### v2

@@ I Introduction
[…]
@@ II-A Parser-based Log Compression
In the realm of log compression, parser-based methods represent the predominant approach. They leverage the inherent semi-structured nature of logs, typically composed of static format strings and dynamic parameters. The process begins with a critical prerequisite step, i.e., log parsing. As illustrated in Fig. 1, parsers like Drain [14] are employed to decompose raw log messages, separating structured headers (e.g., timestamps, log levels) from the free-form body, and then segment the body into an invariant event template and its corresponding variables. This structured representation enables the core compression mechanism, where repetitive template text is replaced with compact identifiers (stored only once in a dictionary) and the extracted parameters are aggregated for subsequent encoding using specialized compression techniques. The resulting output, which is a highly regular stream of template IDs and parameter arrays with significantly reduced entropy, is then processed by general-purpose algorithms like LZMA or gzip to eliminate any remaining statistical redundancy. Several representative log compressors [23, 33, 21, 38] exemplify different design philosophies within this framework.
@@ II-C A Motivating Example for Unified Redundancy Encoding
A straightforward way to implement this idea is to build a prefix tree (i.e., a trie) over the sequences of log tokens and assign IDs to frequent paths, thereby identifying recurring log (sub)sequences. However, this naive construction presents two fundamental limitations. First, high-cardinality fields induce many branches, causing the number of nodes to grow with the product of distinct values per position. This results in prohibitive storage consumption and increased lookup latency due to structural inefficiency. Second, a trie only encodes prefixes. If an infrequent token appears early, the match stops and the rest of the log line is no longer compressible, even when the suffix is highly regular (e.g., uid=509 euid=0 after a rare rhost in L6). These constraints motivate our design of LogPrism, an effective log compressor which moves beyond pure prefix matching and supports compact, gap-tolerant aggregation of frequent token sequences.
@@ III-B Variable Subtree Encoding
The construction of the structural context tree leverages only stable tokens, resulting in a highly compact skeleton where all variable tokens are aggregated at the terminal nodes. To capture the deep correlations between the log structure and its parameters, we can conceptually extend the URT by attaching a prefix subtree to each terminal node, using the varLists of the associated log entries as input paths. By doing so, we unify the structural skeleton and variables into a single, continuous representation. However, this presents two critical challenges. First, a naive sequential insertion of variables inevitably triggers a secondary branch explosion. As illustrated in Fig. 9(a), if variables are processed in their original order (e.g., $v_{0}\to v_{1}\to v_{2}\to v_{3}$), a high-cardinality variable may appear early in the sequence (e.g., a rhost at $v_{1}$), forcing the subtree to branch at the top layers. Consequently, even if later variables form a highly frequent pattern (e.g., uid=509 $\to$ euid=0), it is fragmented into multiple separate branches, preventing efficient compression. Second, we require a robust mechanism to distinguish meaningful patterns from random noise. Since variables are high-entropy, assigning a unique ID to every combination would result in a prohibitively large dictionary and compromise the compression benefits. Therefore, we must selectively encode only frequent, co-occurring variable combinations (the “signal”) into the URT, while actively filtering out rare or random values (the “noise”) to be handled as residuals (Sec. III-C).
@@ III-B1 Variable Reordering for Efficient Subtree Construction
The optimization begins with high-frequency filtering, which prunes infrequent variables, as shown in Fig. 10(a). For each variable position within a log group, we compute the frequency of every unique value and discard any that fall below a pre-defined threshold $\tau$, e.g., the rare rhost=srv2.alfahost.nl. If all values at a given position are filtered out, the entire position is excluded from subsequent analysis. For instance, in the httpd log group (pathID=1) shown in Fig. 8, the sole variable $v_{0}$ (transaction_id) is removed entirely because both of its values (5001 and 5002) are too infrequent. In this work, we use $\tau$ as our universal threshold for different filtering operations to screen out outliers, which avoids the overhead of managing multiple distinct thresholds and reduces overall system complexity.
@@ III-B2 Subtree Construction and Unified Redundancy Encoding
The construction process is illustrated in Fig. 10(b). As each log’s varList is traversed according to the optimal order (i.e., $v_{3}\to v_{0}\to v_{2}\to v_{1}$), a path is extended in the subtree. Every node in this tree maintains a cnt attribute, a key metric that records the number of logs whose reordered variables share that specific prefix. For logs like L5 and L7, which are composed entirely of frequent variables, a complete path is formed, and the cnt value of every node along that path is incremented. In contrast, the traversal for log L6 terminates prematurely because its variable rhost=srv2.alfahost.nl was filtered out as a low-frequency value. Consequently, L6 only increments the cnt values for its matched prefix (euid=0 $\to$ user=test1 $\to$ uid=509). This distinction is crucial, as the cnt attribute now precisely tracks the frequency of both complete and partial patterns. Next, a pruning operation is performed, which acts as a second layer of filtering, removing paths composed of individually frequent variables but whose combination is rare. Any node whose cnt falls below the pruning threshold $\tau$ is removed. While no nodes are pruned in our simplified example (Fig. 10), this is crucial for eliminating noise in complex datasets.
The most critical step is the assignment of new universal pathIDs. To maximize compression efficiency, identifiers are assigned only to nodes that represent the termination of a high-frequency aggregate pattern. This ensures that every pathID corresponds to a meaningful, recurring token combination, preventing an explosion in dictionary size due to rare partial matches. We term these nodes “stable endpoints” and identify them using the cnt attribute under two conditions. First, any leaf node is inherently a stable endpoint, as it represents the explicit termination of a pattern that has already survived the high-frequency filtering process (i.e., $\tau$). For example, the rhost=pokemon1.cs.edu node in Fig. 10(b) is assigned pathID=5 because it marks the end of the pattern for logs L5 and L7. Second, a non-leaf node is designated a stable endpoint if the difference between its cnt and the sum of its children’s cnt values is not less than the threshold $\tau$. This “residual count” represents the number of log entries whose pattern matches exactly up to this node but does not continue to any of its high-frequency children. By enforcing the threshold, we ensure that we only create a new ID if a substantial number of logs terminate at this specific intermediate point. For instance, the uid=509 node has a cnt of 3, while its only child has a cnt of 2. The residual count of 1 indicates that one log (L6) terminates its match here. If this residual meets the threshold, the node is marked as a stable endpoint and assigned pathID=4, allowing LogPrism to hierarchically encode both complete patterns and frequent sub-patterns.
@@ III-C2 Just-in-Time Residual Templatization
Although this stage borrows the principle of numeric classification from Denum, our hierarchical redundancy mining strategy is fundamentally different. Denum performs an indiscriminate, global numeric/non-numeric split from the outset, which can prematurely destroy valuable contextual correlations. In contrast, LogPrism has already processed the vast majority of high-frequency tokens (including many numerics) via unified redundancy encoding in the first two stages. Consequently, the computationally intensive parsing logic of this stage is applied only to a minimal subset of true residual data. This targeted approach provides a dual advantage: a higher compression ratio, as it preserves the “structure + variable” correlations, and faster compression speed, by confining the expensive variable processing to a much smaller dataset.
@@ IV-B RQ1: The Compression Ratio of LogPrism
Table II presents the compression ratio comparison across all benchmarks. LogPrism achieves the highest compression ratio on 14 of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
@@
Comparison with log-specific compressors: LogPrism advances the state-of-the-art among log-specific methods. Compared to LogReducer, CR improvements reach up to 59.26%. Against LogShrink, the maximum gain is 56.63%. Notably, even compared to the current leading method Denum, LogPrism achieves substantial improvements of up to 25.93% on Linux and 18.86% on Thunderbird. The LogZip result for Thunderbird is marked as unavailable (–) because, as reported in the LogShrink paper [21], LogZip failed to complete parsing within one week.
@@
Android 7.742 18.857 12.787 19.370 25.165 20.776 21.857 32.494 32.822
Apache 21.308 25.186 29.557 31.688 30.375 43.028 55.940 58.517 64.330
BGL 12.927 17.637 15.461 18.927 32.655 38.600 42.385 41.804 47.145
Hadoop 20.485 36.095 32.598 32.110 35.008 52.830 60.091 78.546 79.733
HDFS 10.636 13.559 14.059 19.155 26.666 22.634 27.319 25.670 26.466
HealthApp 10.957 13.431 13.843 15.337 12.632 31.694 39.072 44.472 50.477
HPC 11.263 15.076 12.756 14.822 27.208 32.070 35.878 45.275 45.447
Linux 11.232 16.677 14.695 18.508 23.368 25.213 29.252 30.449 38.343
Mac 11.733 22.159 18.074 28.469 26.306 35.251 39.860 40.789 44.992
OpenSSH 16.828 18.918 22.865 31.977 42.606 86.699 103.175 101.654 113.327
OpenStack 12.158 14.437 15.231 17.429 17.258 16.701 22.157 22.238 23.356
Proxifier 15.716 18.982 23.619 25.489 21.493 25.501 27.029 27.288 29.849
Spark 17.825 19.908 26.497 30.614 20.825 59.470 59.739 59.470 54.813
Thunderbird 16.462 27.309 25.428 33.026 49.185 48.434 63.824 75.862
Windows 17.798 202.568 67.533 61.083 310.596 342.975 456.301 481.350 497.298
Zookeeper 25.979 27.667 36.156 38.931 47.373 94.562 116.981 135.251 143.742
This systematic performance advantage of LogPrism stems from a fundamental architectural shift. Traditional methods (whether the “parse-then-compress” workflows or Denum’s “numeric/non-numeric” split) perform an irreversible token categorization early in the process. This leads to the loss of contextual correlations that spans these category boundaries. In contrast, LogPrism’s unified redundancy encoding paradigm co-designs structural extraction with pattern encoding. This enables representing entire, highly correlated “structure + variable” token sequences with single path identifiers, where other methods require one template ID plus multiple independent variable IDs. Our hierarchical redundancy mining strategy prioritizes encoding high-value aggregate patterns while preserving full context, allowing it to exploit deep redundancies in log data.
On the two datasets where LogPrism does not achieve the highest CR (HDFS and Spark), performance remains highly competitive. On HDFS, the difference from optimal is minimal (3.12%). The larger gap on Spark reflects the characteristics of the dataset, i.e., Spark logs are dominated by simple, repetitive templates with single continuously changing numeric values (e.g., over a million records like "Update row <*>"[38]. In this case, the deep correlation mining of LogPrism’s second stage offers less advantage compared to methods specifically optimized for simple numeric streams. Nevertheless, LogPrism achieves a better performance balance in these cases with significantly higher compression speed, as we demonstrate in RQ2.
Summary for RQ1: LogPrism surpasses existing general-purpose and log-specific compressors in compression ratio on 14 out of 16 datasets, establishing a new state-of-the-art in compression effectiveness.
@@ IV-C RQ2: The Compression Speed of LogPrism
The observed variation in LogPrism’s speed across datasets is primarily driven by dataset size, which dictates the effectiveness of chunk-level parallelization. For example, on the small Linux dataset (25,567 lines), the volume is insufficient to fill even a single 100K-line processing chunk, precluding the standard multi-threaded acceleration used by all compressors. Despite this constraint, the standard LogPrism configuration achieves the best performance (8.55 MB/s) among all baselines. Moreover, LogPrism-P provides substantial additional speedup by leveraging the co-design of hierarchical redundancy mining and internal fine-grained parallelism. This ensures high efficiency even in edge cases. On the Linux dataset, LogPrism-P (10.23 MB/s) is significantly faster than Denum (4.97 MB/s) because its hierarchical strategy minimizes the computational overhead, while the internal worker threads maximize CPU utilization within the single active chunk.
@@ IV-D RQ3: Ablation Analysis of LogPrism’s Different Stages
Android 18.857 25.564 32.822
Apache 25.186 61.887 64.330
BGL 17.637 46.548 47.145
Hadoop 36.095 75.170 79.733
HDFS 13.559 24.282 26.466
HealthApp 13.431 46.932 50.477
HPC 15.076 43.880 45.447
Linux 16.677 37.356 38.343
Mac 22.159 37.373 44.992
OpenSSH 18.918 95.969 113.327
OpenStack 14.437 21.868 23.356
Proxifier 18.982 26.664 29.849
Spark 19.908 50.656 54.813
Thunderbird 27.309 59.876 75.862
Windows 202.568 490.787 497.298
Zookeeper 27.667 118.531 143.742
Compression Speed Analysis: The ablation study also reveals the critical role of Stage 2 as a performance accelerator. Despite adding another analysis step, the full LogPrism model is $25.93\%$ faster on average than the structurally simpler LogPrism (S1+S3) variant. This counter-intuitive result validates our hierarchical redundancy mining strategy. In the LogPrism (S1+S3) configuration, every variable identified in Stage 1 must be processed by Stage 3, which relies on computationally expensive regex templatization. In contrast, the full LogPrism model only supplies Stage 3 with a smaller set of “true long-tail residuals.” By handling the majority of high-frequency variables in the efficient Stage 2, LogPrism drastically reduces the workload of the most time-consuming part of the pipeline, thereby lowering overall computational overhead and increasing throughput.